OpenStudy (anonymous):

Ad revenue for a web service is about $2 a month per each free user. A monthly “premium” subscription costs $11 dollars for the month, but does not make any ad revenue. The first month showed that 690 users used the web service that month earning a total of $2,064 between ad revenue and subscriptions. Determine the system of equations describing this relationship. Calculate the number of users of each type(free users,subscribers)

OpenStudy (anonymous):

Anybody got any idea?

OpenStudy (here_to_help15):

i might :P

OpenStudy (anonymous):

:DD

OpenStudy (here_to_help15):

let there be 'f' free users and 'p' premium users total users = f+p given: total users for first month = 690 so we have : f + p = 690 => p = 690 - f revenue 1 free user = $2 revenue by 'f' free users = $2*f revenue by 1 premium user = $11 revenue by 'p' premium user = $11*p total revenue = (2*f) + (11*p) = $2064 => 2f + 11(690-f) = 2064 solve for f. then use it to get the value of p.

OpenStudy (anonymous):

I remember learning something like this maybe 2 years ago but naw I'm soz i don't remember lol

OpenStudy (here_to_help15):

what grde is this again

OpenStudy (here_to_help15):

i remeber taking it very currently

OpenStudy (anonymous):

12

OpenStudy (here_to_help15):

yup see told u so took it very current

OpenStudy (here_to_help15):

and theres ur idea

OpenStudy (anonymous):

i remember learning something like it but who payed attention in high school lol

OpenStudy (here_to_help15):

i did :P :) just in case i needed it someday

OpenStudy (anonymous):

ehh i thought it was useless i couldn't wait for college 2 where i would learn things that would contribute to what i wanted to do in life lol

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

so where u at now in life

OpenStudy (anonymous):

College XD going to school for car design

OpenStudy (here_to_help15):

oh cool :)

OpenStudy (anonymous):

yea its ight i wanted to be a negotiator but the field is so limited

OpenStudy (anonymous):

Well if you ever got any history questions lemme know i love history

OpenStudy (here_to_help15):

i was going for secret service lol

OpenStudy (here_to_help15):

or something in that category

OpenStudy (anonymous):

What made you want to work with those guys lol

OpenStudy (here_to_help15):

a lot lol

OpenStudy (anonymous):

You going to school for it or like?

OpenStudy (here_to_help15):

yea need bachelor

OpenStudy (anonymous):

ahh ok

OpenStudy (here_to_help15):

yup :D

OpenStudy (anonymous):

seems like a limited field lol obama has enough protection as it is XD

OpenStudy (here_to_help15):

LOL

OpenStudy (here_to_help15):

XD good 1 lol

OpenStudy (here_to_help15):

well u dont just protect the president humph lol

OpenStudy (anonymous):

With his little trespassers running on his lawn and crap, like the main new coverage.

OpenStudy (anonymous):

news*

OpenStudy (here_to_help15):

hahaha

OpenStudy (here_to_help15):

mk tht or FBI

OpenStudy (here_to_help15):

both need bachelor

OpenStudy (anonymous):

While you're looking over crime cases I'm going to be drawing maseratis lol

OpenStudy (here_to_help15):

lol

OpenStudy (anonymous):

ill do anything but be an accountant like my father lol

OpenStudy (here_to_help15):

lol

OpenStudy (anonymous):

i wonder what julie here plans on doing considering she's in 12th grade lol

OpenStudy (here_to_help15):

yeah lol

OpenStudy (here_to_help15):

hby julie got any plans?

OpenStudy (anonymous):

Guess not :3

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

yeah more secret service then FBI